Treasury bills are currently paying 6 percent and the inflation rate is 3.10 percent.
 
What is the approximate real rate of interest? 
 
What is the exact real rate? 
Answer #1

approximate real rate = nominal - iflation = 6-3.1 = 2.9%

real rate = (nominal-inflation)/(1+inflation) = (.06-.031)/(1.031) = 2.81%